OpenStudy (anonymous):

Help! Ivan walks to school one mile/hr. faster than when he walks back home. If it takes him 12 min. (0.20 hr.) to walk to school and 18 min. (0.30 hr.) to walk back, how fast does he walk to school and how far does he have to walk to school?

OpenStudy (anonymous):

@PhiBeeandJay u can use the simply formula: v = s/t where v is speed, s is distance or displacement, and t is time

OpenStudy (anonymous):

@PhiBeeandJay u got it??

OpenStudy (anonymous):

please try

OpenStudy (anonymous):

I'm trying to come up with equation you just gave me.

OpenStudy (anonymous):

Distance is your constant value so you can set the to and from school distances equal to each other: \[s_1=s_2\] use the equation that @gerryliyana gave you but in a different form so that you can get rid of the distance. You will also have to introduce a variable x is the speed that Ivan walks home from school: \[s_1*t_1=s_2*t_2\] \[(x+1)*0.2=x*0.3\]

OpenStudy (anonymous):

vs = 1 mil/hr + vh s/0.2 = 1+ (s/0.3) So s?? (where vs is speed Ivan walks to school , and vh is speed he walks back home )

OpenStudy (anonymous):

do I multiply by 0.2 on both sides now?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

@PhiBeeandJay : So What Did You Get?